# On the Convergence of Multi-objective Optimization under Generalized Smoothness

### v3

@@ 3.2 Stochastic Generalized Smooth Multi-objective Gradient descent (SGSMGrad)
2:  $w_{0}$=Warm-start($w_{0}$, $x_{0}$, $\rho$) for small-level iteration-wise CA distance only
3:  for $t=0,1,...,T-1$ do
4:     $x_{t+1}=x_{t}-\alpha\nabla G_{1}(x_{t})w_{t}$
5:     $w_{t+1}=\Pi_{\mathcal{W}}\big{(}w_{t}-\beta[\nabla G_{2}(x_{t})^{\top}\nabla G_{3}(x_{t})w_{t}+\rho w_{t}]\big{)}$
6:  end for
3.3 Fast approximation via Taylor expansion
Similarly to most MGDA-type algorithms, our methods require $\mathcal{O}(K)$ space and time to compute and store all task gradients at each iteration for updating the weight $w_{t}$. This becomes a drawback when the number of tasks or the model size is large. Motivated by [23], one solution is to use the Taylor Theorem to approximate the gradient for updating the weight $w_{t}$ as
@@ 3.3 Fast approximation via Taylor expansion
where $R(x_{t})$ is the remainder term and it takes the order $R(x_{t})=o(\|x_{t}-x_{t+1}\|^{2})$, which can be made sufficiently small by adjusting the step size. Thus, we then propose GSMGrad-FA in Algorithm 4 (shown in Appendix B), where we update $x_{t}$ along the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to get $x_{t+1}$ following by the update rule of $w_{t}$
@@
As a result, in the model parameters update process, we only require one backward process by calculating the gradient of $F(x_{t})w_{t}$ w.r.t. $x_{t}$ without storing it, and additional forward processes to compute $F(x_{t+1})$ in the weight update process. This approach saves computational and memory costs in the practical implementation significantly. More importantly, we also provide a theoretical guarantee for this efficient method (in eq. 6).
@@ 4.1 Deterministic setting
Each objective function $f_{i}$ $\forall i\in[K]$ is $\ell$-smooth defined in Definition 1, where :[0,+)($\ell:[0,+\infty)\rightarrow(0,+\infty)$ is a continuous non-decreasing function such that $\varphi(a)=\frac{a^{2}}{2\ell(2a)}$ is monotonically increasing for any $a\geq 0$.
@@
Let Assumptions 1 and 2 hold. Set $\beta\sim\mathcal{O}(\frac{1}{M^{2}}),\alpha\sim\mathcal{O}(\frac{1}{M^{2}}+\frac{1}{M\ell(M+1)}),T\geq\max\Big{(}\Theta\big{(}\frac{1}{\alpha\epsilon^{2}}\big{)},\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\Big{)}$ and $\rho\sim\mathcal{O}(\epsilon^{2})$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and detailed proof can be found in Appendix C.1. Theorem 1 provides the first convergence rate to obtain an $\epsilon$-accurate Pareto stationary point for MOO problems with $\ell$-smooth objectives. Moreover, it achieves the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-2})$for GD with a single standard $L$-smooth objective[5]. The MOO problems with $\ell$-smooth objectives are challenging due to two reasons: 1) $\|\nabla F(x)\|$ is potentially unbounded in our $\ell$-smoothness setting, making all existing analysis in MOO [24, 13, 27, 32, 6, 33, 3] not applicable. 2) the update of $x$ includes all gradient information from each task, making the existing adaptive methods for single generalized smooth functions invalid.
@@
Under the same setting in Theorem 1, $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}=\mathcal{O}(\epsilon^{2})$.
@@ 4.2 Stochastic setting
Let $F,c>0$ and $0<\delta\leq\frac{1}{2}$ be some constants such that $F\geq\frac{4(\Delta+c)}{\delta}$ and $M=\sup\{z\geq 0|\varphi(z)\leq F\}$. Define the following random variables $\tau_{1}=\min\{t|\exists i\in[K],f_{i}(x_{t+1})-f_{i}^{*}>F\}\wedge T,\tau_{2}=\min\{t|\exists i\in[K],j\in[3],\|\varepsilon_{t,j,i}\|>\frac{L_{0}}{\sqrt{\alpha\rho}}\}\wedge T,\tau_{3}=\min\{t|\exists i,j\in[K],\|\varepsilon_{t,2,i}\|\|\varepsilon_{t,3,j}\|>\frac{L_{1}}{\sqrt{\alpha\rho}}\}\wedge T$ and $\tau=\min\{\tau_{1},\tau_{2},\tau_{3}\}$, where $L_{0},L_{1}>0$ are some constants and $a\wedge b$ denotes $\min(a,b)$. We have the following theorem:
@@
Let Assumptions 1, 2, and 3 hold. Set $\alpha\leq\min\{\mathcal{O}(\ell(M+1)),\mathcal{O}(\beta),\mathcal{O}(\rho),\mathcal{O}\left(\frac{1}{\sqrt{T}}\right),\mathcal{O}(\epsilon^{2})\}$, $\beta\leq\min\{\mathcal{O}\left(\frac{1}{\alpha{T}}\right),\mathcal{O}(\epsilon^{2}),\mathcal{O}(\rho)\}$, $\rho\leq\min\{\mathcal{O}\left(\frac{1}{\alpha T}\right),\mathcal{O}(\frac{1}{\sqrt{\alpha T}}),\mathcal{O}\left(\frac{\epsilon}{\sqrt{\beta}}\right),\mathcal{O}(\epsilon^{2})\}$ and $T\geq\Theta\left(\frac{1}{\alpha\epsilon^{2}}+\frac{1}{\epsilon^{4}}\right)$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2},$ with the probability at least $1-\delta$.
The full version and detailed proof can be found in Appendix C.3. When we set $\alpha,\beta,\rho\sim\mathcal{O}(\epsilon^{2})$ and $T\sim\mathcal{O}(\epsilon^{-4})$, we can find an $\epsilon$-stationary point with the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-4})$ for SGD with a single $L$-smooth objective[1]. Note that in the proof of Theorem 1, we show for each $t\leq T$ and $w\in W$, we have that $F(x_{t})w$ is bounded by applying a small constant step size $\alpha$ and $\beta$. However, this condition does not necessarily hold for our stochastic setting due to the unbounded gradient noise. To solve this problem, we introduce stopping time $\tau$. The advantages are as follows: 1) for any $t\leq\tau$, $w\in W$, we have that $F(x_{t})w$ is bounded; 2) for any $t<\tau$, the norm of gradient noise is bounded; 3) due to the optimal stopping theorem, for any $w\in W$ and $i\in[3]$, we have that $\mathbb{E}[\sum_{t=0}^{\tau}\varepsilon_{t,i}w]=0$. Based on these properties, we can further get the following lemma:
@@
The proof of Lemma 1 is available in C.4. Lemma 1 indicates that $\frac{\alpha}{2}\mathbb{E}\left[\sum_{t=0}^{\tau-1}\|\nabla F(x_{t})w_{t}\|^{2}\right]$ is bounded by some constant and if $\tau=T$ with high probability, we have that $\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\Big{|}\tau=T\right]\sim\mathcal{O}\left(\frac{1}{\alpha T}\right)$. Note that $\{\tau<T\}=\{\tau_{2}<T\}\cup\{\tau_{3}<T\}\cup\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\}.$ The first two events are related to the gradient noise, where the probabilities can be bounded by Assumption 3 and Chebyshev’s inequality. The last event indicates that for some $i\in[K]$, we have $f_{i}(x_{\tau})-f_{i}^{*}\leq\frac{F}{2}.$ Based on Lemma 1 and Markov inequality, we can show that ($\mathbb{P}(\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\})\leq\frac{\delta}{4}$ and we can further show that $\mathbb{P}(\tau=T)\geq 1-\frac{\delta}{2}$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}$ converges with high probability. Similar to Corollary 1, Theorem 2 also implies the average of CA distances converges with time with high probability.
@@ 5 Convergence Analysis under Iteration-wise CA distance
In Section 4 we show that the average CA distance is bounded under generalized smooth conditions. The average CA distance is also studied in MoCo [13] and MoDo [6] with the bounded gradient assumption and these works only focus on guarantees of the average CA distance over iterations. However, a $\epsilon$-level average CA distance only implies the smallest CA distance to be $\epsilon$-level. Since we want to keep the update direction close enough to the CA direction, it is better to have a tighter bound of CA distances. In this section, we show the CA distance is $\mathcal{O}(\epsilon)$ at every iteration with the help of a warm-start process and convergence results for Algorithms 1, 3, and 4.
@@ 5.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}},\rho\sim\mathcal{O}(\epsilon^{2}),\beta\sim\mathcal{O}(\epsilon^{4}),\alpha\sim\mathcal{O}(\epsilon^{9})$, $N\sim O(\epsilon^{-2})$ as constants, and $T\sim\Theta(\epsilon^{-11})$. All the parameters satisfy the requirements in the formal version of Theorem 1 and we have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\sim\mathcal{O}(\epsilon).$
The finite time error bound and the full proof can be found in Appendix D.1. Since our parameters satisfy the requirements in the formal version of Theorem 1, we can find an $\epsilon$-accurate Pareto stationary point with $\mathcal{O}(\epsilon^{11})$ samples. In the analysis of CA distance, we show that the CA distance can be bounded by the term $\|w_{t}-w_{t,\rho}^{*}\|$ plus the strongly-convex constant $\rho$. Meanwhile, there is a decay relation between $\|w_{t+1}-w_{t+1,\rho}^{*}\|$ and $\|w_{t}-w_{t,\rho}^{*}\|$ with some error terms controlled by step sizes. Nevertheless, the error terms will accumulate since we do telescoping on this decay relation, which will be the dominating term. Thus, step sizes have to be much smaller than the choices in Theorem 1 to guarantee iteration-wise small CA distance.
@@
Let Assumptions 1 and 2 hold. Set $N\sim O(\epsilon^{-2})$, $\beta^{\prime}\leq\frac{1}{M^{2}},\rho\leq\min\left(\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\alpha T}\right)\right),\beta\leq\mathcal{O}(\epsilon^{2}),\alpha\leq\min\left(\mathcal{O}(\beta),\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\beta T}\right)\right)$ as constants, $T\geq\max\left(\Theta\left(\frac{1}{\alpha\epsilon^{2}}\right),\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\right)$. We have $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and proof can be found in the Section D.2. We can easily extend the analysis in Section C.1 on convergence analysis of Algorithm 4 because the only extra effort is dealing with the remainder term, which can be simply bounded by the smallest step size. As a result, the sample complexity remains the same $\mathcal{O}(\epsilon^{-11})$ to achieve a Pareto stationary point.
@@
Let Assumptions 1 and 2 hold. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},\rho\sim\mathcal{O}(\epsilon^{2}),\beta\sim\mathcal{O}(\epsilon^{4}),\alpha\sim\mathcal{O}(\epsilon^{9})$, $N\sim O(\epsilon^{-2})$ as constants, and $T\sim\Theta(\epsilon^{-11})$. We have that $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\sim\mathcal{O}(\epsilon).$

### v4

@@ 3.2 Stochastic MGDA with Double Sampling
2:  Option I: $w_{0}$=warm-start($w_{0}$, $x_{0}$, $\rho$) # for analyzing iteration-wise CA distance
3:  Option II: $w_{0}\leftarrow w_{0}$                     # for analyzing averaged CA distance
4:  for $t=0,1,...,T-1$ do
5:     $x_{t+1}=x_{t}-\alpha\nabla G_{1}(x_{t})w_{t}$
6:     $w_{t+1}=\Pi_{\mathcal{W}}\big{(}w_{t}-\beta[\nabla G_{2}(x_{t})^{\top}\nabla G_{3}(x_{t})w_{t}+\rho w_{t}]\big{)}$ double sampling
7:  end for
3.3 MGDA with Fast Approximation
It can be seen from Algorithm 1 and Algorithm 3 that MGDA requires $\mathcal{O}(K)$ space and time to compute and store all task gradients at each iteration for updating the weight $w_{t}$. This becomes a drawback when the number of tasks or the model size is large. Motivated by [24], one solution is to use the Taylor Theorem to approximate the gradient for updating the weight $w_{t}$ as
@@ 3.3 MGDA with Fast Approximation
where $R(x_{t})$ is the remainder term and it takes the order $R(x_{t})=o(\|x_{t}-x_{t+1}\|^{2})$, which can be made sufficiently small by adjusting the step size. By incorporating this fast approximation (FA) in Algorithm 1, we then present MGDA-FA in Algorithm 4 (shown in Appendix B), where $x_{t}$ is updated along the update direction $d_{t}=\nabla F(x_{t})w_{t}$ to get $x_{t+1}$ following by the update rule of $w_{t}$
@@
As a result, in the model parameters update process, MGDA-FA only requires one backward process by calculating the gradient of $F(x_{t})w_{t}$ w.r.t. $x_{t}$ without storing it, and additional forward processes to compute $F(x_{t+1})$ in the weight update process. This approach saves computational and memory costs in the practical implementation significantly. Most importantly, we provide a theoretical guarantee for this efficient method (in eq. 6).
@@ 4.1 Deterministic setting
Each objective function $f_{i}$ $\forall i\in[K]$ is generalized $\ell$-smooth defined in Definition 1, where :[0,+)($\ell:[0,+\infty)\rightarrow(0,+\infty)$ is a continuous non-decreasing function such that $\varphi(a)=\frac{a^{2}}{2\ell(2a)}$ is monotonically increasing for any $a\geq 0$.
@@
Let Assumptions 1 and 2 hold. Set $\beta\leq\mathcal{O}(\frac{1}{M^{2}}),\alpha\leq\mathcal{O}(\frac{1}{M^{2}}+\frac{1}{M\ell(M+1)}),T\geq\max\Big{(}\Theta\big{(}\frac{1}{\alpha\epsilon^{2}}\big{)},\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\Big{)}$ and $\rho=\mathcal{O}(\epsilon^{2})$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and detailed proof can be found in Appendix C.1. Theorem 1 provides the first convergence rate to obtain an $\epsilon$-accurate Pareto stationary point for MOO problems with generalized $\ell$-smooth objectives. Moreover, it achieves the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-2})$for GD with a single standard $L$-smooth objective [5]. The MOO problems with generalized $\ell$-smooth objectives are challenging due to two reasons: 1) $\|\nabla F(x)\|$ is potentially unbounded in our generalized $\ell$-smoothness setting, making all existing analysis in MOO [25, 13, 29, 35, 6, 36, 3] not applicable. 2) the update of $x$ includes all gradient information from each task, making the existing adaptive methods for single generalized smooth functions invalid.
@@
Under the same setting in Theorem 1, $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|^{2}\leq\mathcal{O}(\epsilon^{2})$.
@@ 4.2 Stochastic setting
Let $F,c>0$ and $0<\delta\leq\frac{1}{2}$ be some constants such that $F\geq\frac{4(\Delta+c)}{\delta}$ and $M=\sup\{z\geq 0|\varphi(z)\leq F\}$. Define the following random variables $\tau_{1}=\min\{t|\exists i\in[K],f_{i}(x_{t+1})-f_{i}^{*}>F\}\wedge T,\tau_{2}=\min\{t|\exists i\in[K],j\in[3],\|\varepsilon_{t,j,i}\|>\frac{L_{0}}{\sqrt{\alpha\rho}}\}\wedge T,\tau_{3}=\min\{t|\exists i,j\in[K],\|\varepsilon_{t,2,i}\|\|\varepsilon_{t,3,j}\|>\frac{L_{1}}{\sqrt{\alpha\rho}}\}\wedge T$ and $\tau=\min\{\tau_{1},\tau_{2},\tau_{3}\}$, where $L_{0},L_{1}>0$ are some constants and $a\wedge b$ denotes $\min(a,b)$. We have the following theorem for Algorithm 3::
@@
Let Assumptions 1, 2, 3 hold. Set $\alpha\leq\min\{\mathcal{O}(\ell(M+1)),\mathcal{O}(\beta),\mathcal{O}(\rho),\mathcal{O}\left(\frac{1}{\sqrt{T}}\right),\mathcal{O}(\epsilon^{2})\}$, $\beta\leq\min\{\mathcal{O}\left(\frac{1}{\alpha{T}}\right),\mathcal{O}(\epsilon^{2}),\mathcal{O}(\rho)\}$, $\rho=\min\{\mathcal{O}\left(\frac{1}{\alpha T}\right),\mathcal{O}(\frac{1}{\sqrt{\alpha T}}),\mathcal{O}\left(\frac{\epsilon}{\sqrt{\beta}}\right),\mathcal{O}(\epsilon^{2})\}$ and $T\geq\Theta\left(\frac{1}{\alpha\epsilon^{2}}+\frac{1}{\epsilon^{4}}\right)$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2},$ with the probability at least $1-\delta$.
The full version and detailed proof can be found in Appendix C.3. When we set $\alpha,\beta,\rho=\mathcal{O}(\epsilon^{2})$ and $T\geq\mathcal{O}(\epsilon^{-4})$, we can find an $\epsilon$-stationary point with the optimal sample complexity in the order of $\mathcal{O}(\epsilon^{-4})$ for SGD with a single $L$-smooth objective [1]. Note that in the proof of Theorem 1, we show for each $t\leq T$ and $w\in W$, we have that $F(x_{t})w$ is bounded by applying a small constant step size $\alpha$ and $\beta$. However, this condition does not necessarily hold for our stochastic setting due to the unbounded gradient noise. To solve this problem, we introduce stopping time $\tau$. The advantages are as follows: 1) for any $t\leq\tau$, $w\in W$, we have that $F(x_{t})w$ is bounded; 2) for any $t<\tau$, the norm of gradient noise is bounded; 3) due to the optimal stopping theorem, for any $w\in W$ and $i\in[3]$, we have that $\mathbb{E}[\sum_{t=0}^{\tau}\varepsilon_{t,i}w]=0$. Based on these properties, we can further get the following lemma:
@@
The proof of Lemma 1 is available in C.4. Lemma 1 indicates that $\frac{\alpha}{2}\mathbb{E}\left[\sum_{t=0}^{\tau-1}\|\nabla F(x_{t})w_{t}\|^{2}\right]$ is bounded by some constant and if $\tau=T$ with high probability, we have that $\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\Big{|}\tau=T\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)$. Note that $\{\tau<T\}=\{\tau_{2}<T\}\cup\{\tau_{3}<T\}\cup\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\}.$ The first two events are related to the gradient noise, where the probabilities can be bounded by Assumption 3 and Chebyshev’s inequality. The last event indicates that for some $i\in[K]$, we have $f_{i}(x_{\tau})-f_{i}^{*}\leq\frac{F}{2}.$ Based on Lemma 1 and Markov inequality, we can show that ($\mathbb{P}(\{\tau_{1}<T,\tau_{2}=T,\tau_{3}=T\})\leq\frac{\delta}{4}$ and we can further show that $\mathbb{P}(\tau=T)\geq 1-\frac{\delta}{2}$. We then have that $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}$ converges with high probability. Similar to Corollary 1, Theorem 2 also implies the average of CA distances converges with time with high probability.
@@ 5 Convergence Analysis under Iteration-wise CA distance
In Section 4 we show that the average CA distance is bounded under generalized smooth conditions. The average CA distance is also studied in MoCo [13] and MoDo [6] with the bounded gradient assumption and these works only focus on guarantees of the average CA distance over iterations. However, a $\epsilon$-level average CA distance only implies the smallest CA distance to be $\epsilon$-level. Since we want to keep the update direction close enough to the CA direction, it is better to have a tighter bound of CA distances. In this section, we show the iterative CA distance is $\mathcal{O}(\epsilon)$ with a warm-start process and convergence results for Algorithms 1, 3, and 4.
@@ 5.1 Deterministic setting
Let Assumptions 1 and 2 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}(\epsilon^{2}),\beta\leq\mathcal{O}(\epsilon^{4}),\alpha\leq\mathcal{O}(\epsilon^{9})$, $N\geq O(\epsilon^{-2})$ as constants, and $T\geq\Theta(\epsilon^{-11})$. All the parameters satisfy the requirements in the formal version of Theorem 1 and we have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\leq\mathcal{O}(\epsilon).$
The finite time error bound and the full proof can be found in Appendix D.1. Since our parameters satisfy the requirements in the formal version of Theorem 1, we can find an $\epsilon$-accurate Pareto stationary point with $\mathcal{O}(\epsilon^{-11})$ samples. In the analysis of CA distance, we show that the CA distance can be bounded by the term $\|w_{t}-w_{t,\rho}^{*}\|$ plus the strongly-convex constant $\rho$. Meanwhile, there is a decay relation between $\|w_{t+1}-w_{t+1,\rho}^{*}\|$ and $\|w_{t}-w_{t,\rho}^{*}\|$ with some error terms controlled by step sizes. Nevertheless, the error terms will accumulate since we do telescoping on this decay relation, which will be the dominating term. Thus, step sizes have to be much smaller than the choices in Theorem 1 to guarantee iteration-wise small CA distance.
@@
Let Assumptions 1 and 2 hold. Set $N\geq O(\epsilon^{-2})$, $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\min\left(\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\alpha T}\right)\right),\beta\leq\mathcal{O}(\epsilon^{2}),\alpha\leq\min\left(\mathcal{O}(\beta),\mathcal{O}(\epsilon^{2}),\mathcal{O}\left(\frac{1}{\beta T}\right)\right)$ as constants, $T\geq\max\left(\Theta\left(\frac{1}{\alpha\epsilon^{2}}\right),\Theta\left(\frac{1}{\beta\epsilon^{2}}\right)\right)$. We have $\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
The full version with detailed constants and proof can be found in the Section D.2. We can easily extend the analysis in Section C.1 to the convergence analysis of Algorithm 4 under the average CA distance, because the only extra effort is dealing with the remainder term, which can be simply bounded by the smallest step size. As a result, the sample complexity remains the same $\mathcal{O}(\epsilon^{-11})$ to achieve a Pareto stationary point.
@@
Let Assumptions 1 and 2 hold. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},\rho=\mathcal{O}(\epsilon^{2}),\beta\leq\mathcal{O}(\epsilon^{4}),\alpha\leq\mathcal{O}(\epsilon^{9})$, $N\sim O(\epsilon^{-2})$ as constants, and $T\geq\Theta(\epsilon^{-11})$. We have that $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\leq\mathcal{O}(\epsilon).$